OpenStudy (anonymous):

A high school science classroom is being remodeled. The original area of the classroom was 725 square feet. The area of the remodeled classroom will be 841 square feet. By how many square feet will the area of the classroom increase? What is the percent of increase? Show your work.

OpenStudy (anonymous):

part 1=841-725=116 part 2=116/725=.16=16%

OpenStudy (anonymous):

you answerd your own question?

OpenStudy (anonymous):

yep..

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

any way is it correct???

OpenStudy (anonymous):

ya